Three Engineering students are selected at random, and each is categorized as requiring Applied Statistics & Data Analysis SS2143 (S) or Applied Probability & Statistics SS2141 (P). Equal proportions of students take these courses. If X-the number of students among the three who require SS2143, the total list of outcomes in S and its associated X values are given by Outcome SPP PSP PPS PSS SPS What is the probability of observing more than 1 student who requires S$2143? A. 1/2 B. 1/3 C. 7/8
Answer #1

P(observing more then 1 student who require SS2143 )=P(having 4 out of 8 equally probable outcome) =4/8 =1/2

option A is correct